OpenStudy (anonymous):

A square has four equal sides. If the perimeter (sum of all the sides) of the square is p, what is the length of each side?

OpenStudy (anonymous):

p/4

OpenStudy (help!!!!):

give me a sec 2 look at this

OpenStudy (anonymous):

1/4 p + 1/4 p + 1/4 p + 1/4 p = 4/4 p = 1p = p

OpenStudy (anonymous):

got it thanks!

OpenStudy (anonymous):

yeah no problem :)